### v1

@@ 3 Regularity in classical existence theory
Assume that, for every $i\in I$, the constraint map ${\bf X}_{i}:X_{-i}^{\mathrm{ad}}\to 2^{X_{i}^{\mathrm{ad}}}$ is lower semicontinuous and that Assumption 3.2 holds. Then the generalized Nash equilibrium problem $G=({\bf X},\mathcal{J})$ admits at least one equilibrium.
Proof.
Equilibria of $G$ correspond to fixed points of the best response map
${\bf{B}}(x)=\prod_{i=1}^{N}{\bf{B}}_{i}(x_{-i}),\qquad{\bf{B}}_{i}(x_{-i})=\operatorname*{argmin}_{x_{i}\in{\bf X}_{i}(x_{-i})}\mathcal{J}_{i}(x_{i},x_{-i}).$
By Assumption 3.1(ii), each ${\bf X}_{i}$ is closed-valued and the set ${\bf X}_{i}(x_{-i})$ is nonempty, convex, and compact for all $x_{-i}\in X_{-i}$. Assumption 3.2 ensures that every ${\bf{B}}_{i}(x_{-i})$ is nonempty; hence ${\bf{B}}(x)$ is well defined and nonempty-valued.
To invoke Kakutani’s fixed-point theorem, we must verify that ${\bf{B}}:X^{\mathrm{ad}}\to 2^{X^{\mathrm{ad}}}$ is upper semicontinuous and has nonempty, convex, and compact values. The latter properties follow directly from the assumptions above. We therefore focus on upper semicontinuity.
By [beer_topologies_1993, Theorem 6.2.5], $\bf{B}$} is upper semicontinuous if and only if, for every closed set $C\subset X^{\mathrm{ad}}$, the inverse image ${\bf{B}}^{-1}(C)=\{x\in X^{\mathrm{ad}}:{\bf{B}}(x)\cap C\neq\emptyset\}$ is closed.
Let $(x^{n})\subset X^{\mathrm{ad}}$ satisfy $x^{n}\to x$ and ${\bf{B}}(x^{n})\cap C\neq\emptyset$ for all $n$. Choose $y^{n}\in{\bf{B}}(x^{n})\cap C$. Since $C$ is compact, a subsequence (not relabeled) satisfies $y^{n}\to y\in C$. Closedness of the graph of $\bf X$} implies $y\in{\bf X}(x)$. To show $y\in{\bf{B}}(x)$, we must verify that each $y_{i}$ minimizes $\mathcal{J}_{i}(\cdot,x_{-i})$ over ${\bf X}_{i}(x_{-i})$. This is where the lower semicontinuity of ${\bf X}_{i}$ is essential.
Indeed, since $x^{n}\to x$, for every $w_{i}\in{\bf X}_{i}(x_{-i})$ there exists a sequence $w_{i}^{n}\in{\bf X}_{i}(x_{-i}^{n})$ with $w_{i}^{n}\to w_{i}$. Then, for each $i\in I$,
$\mathcal{J}_{i}(y_{i},x_{-i})=\lim_{n\to\infty}\mathcal{J}_{i}(y_{i}^{n},x_{-i}^{n})\leqslant\lim_{n\to\infty}\mathcal{J}_{i}(w_{i}^{n},x_{-i}^{n})=\mathcal{J}_{i}(w_{i},x_{-i}),$
where the equalities follow from continuity of $\mathcal{J}_{i}$ and the inequality holds because $y_{i}^{n}$ minimizes $\mathcal{J}_{i}(\cdot,x_{-i}^{n})$ over ${\bf X}_{i}(x_{-i}^{n})$. Hence $y_{i}\in{\bf{B}}_{i}(x_{-i})$ for all $i$, so $y\in{\bf{B}}(x)$. Consequently, ${\bf{B}}^{-1}(C)$ is closed, proving that $\bf{B}$} is upper semicontinuous.
By Theorem 3.3, $\bf{B}$} admits a fixed point, which is an equilibrium of $G$. ∎
The argument above makes clear how the lower semicontinuity of ${\bf X}_{i}$ enters the proof and why it is crucial for the closedness of the best-response graph. While standard in abstract fixed-point theory, this assumption is rarely easy to verify in practical settings such as PDE-constrained or function space games, see [bensoussan_impulse_1987, hintermuller_recent_2019, facchinei_generalized_2007, facchinei_decomposition_2011]. This limitation partly explains the popularity of models where the coupling constraints have the structured form (2.2), for which the lower semicontinuity issue can be bypassed by exploiting properties of the Nikaido-Isoda function, as discussed next.
4 Structure versus regularity in the existence theory
This section serves as a bridge between the classical Kakutani approach and the structural conditions in Section 5. We first recall the Nikaido–Isoda framework and the notion of variational equilibrium in jointly convex and potential games. We then explain how these structural assumptions can be viewed as ways to replace lower semicontinuity of the constraint maps by convexity and potential-type properties.
In this section, we clarify how structural assumptions (such as convexity, shared constraints, or potential structure) can be traded against regularity of the constraint maps in existence results for convex GNEPs.
4.1 Jointly convex games, and potential games
The Nikaido-Isoda function $\Psi:X\times X\to\mathbb{R}$ is defined as
@@ 4.1 Jointly convex games, and potential games
Under Assumption 3.2, $\Psi$ is continuous and, for each $x\in X$, the map $\Psi(x,\cdot)$ is convex. The next lemma is central to the state of the art in the theory of GNEPs, therefore we present a short proof.
Lemma 4.1.
A strategy bundle $\overline{x}\in X^{\rm{ad}}$ is a Nash equilibrium of the game $G=({\bf X},\mathcal{J})$ if and only if $\overline{x}\in{\bf X}(\overline{x})$ and $\Psi(\overline{x},\overline{x})\leqslant\Psi(\overline{x},y)$, for all $y\in{\bf X}(\overline{x})$.
Proof.
Let $\overline{x}\in X^{\rm{ad}}$ be a Nash equilibrium. By definition of equilibrium, $\overline{x}$ is a fixed point of $\bf X$}. Moreover, for each $i$ and each $y_{i}\in{\bf X}_{i}(\overline{x}_{-i})$ we have $\mathcal{J}_{i}(\overline{x})\leqslant\mathcal{J}_{i}(y_{i},\overline{x}_{-i}),$ and summing over $i$ yields $\Psi(\overline{x},\overline{x})\leqslant\Psi(\overline{x},y)$ for all $y\in{\bf X}(\overline{x}).$
Conversely, assume that $\overline{x}$ is a fixed point of $\bf X$} and that $\Psi(\overline{x},\cdot)$ reaches its global minimum at $\overline{x}.$ If $\overline{x}$ is not a Nash equilibrium, then for some $i\in I$, there exists $y_{i}\in{\bf X}_{i}(\overline{x}_{-i})$ such that $\mathcal{J}_{i}(\overline{x})>\mathcal{J}_{i}(y_{i},\overline{x}_{-i})$. Noticing that $(y_{i},\overline{x}_{-i})\in{\bf X}(\overline{x})$, the result follows because
@@
In other words, $\overline{x}$ is a GNE if and only if it solves the minimization problem
$\min\Psi(\overline{x},y)\ \ {\rm s.t.}\ \ y\in{\bf X}(\overline{x}).$ (4.2)
@@
Among other consequences, the above lemma allows one to characterize Nash equilibria – under extra regularity of the objective functions – as the solution of a quasi-variational inequality (QVI).
Lemma 4.2.
Assume that, for each $i\in I$ and each $x_{-i}\in X_{-i}$, the function $\mathcal{J}_{i}(\cdot,x_{-i})$ is continuously Gâteaux-differentiable with bounded derivative $\partial_{i}\mathcal{J}_{i}(\cdot,x_{-i}):X_{i}\to X_{i}^{*}.$ Hence $\overline{x}$ is a GNE if, and only if, $\overline{x}\in{\bf X}(\overline{x})$ and
for all $y\in{\bf X}(\overline{x}).$
Remark 4.3.
The differentiability assumption in Lemma 4.2 can be relaxed. If each $\mathcal{J}_{i}(\cdot\!,x_{-i})$ is convex and lower semicontinuous, the equilibrium condition remains valid in the following subdifferential form: $\overline{x}\in{\bf X}(\overline{x})$ is a GNE if and only if
$\sum_{i=1}^{N}\langle\xi_{i},y_{i}-\overline{x}_{i}\rangle_{X_{i}^{*},X_{i}}\geqslant 0,\qquad\text{for all }y\in{\bf X}(\overline{x}),$
for some $\xi_{i}\in\partial_{i}\mathcal{J}_{i}(\overline{x})$, where $\partial_{i}$ denotes the convex subdifferential with respect to $x_{i}$. When $\mathcal{J}_{i}(\cdot,x_{-i})$ is continuously Gâteaux-differentiable, $\partial_{i}\mathcal{J}_{i}(\overline{x})=\{\partial_{i}\mathcal{J}_{i}(\overline{x})\}$ and the statement reduces to (4.3).
Under the assumptions of Theorem 3.4, the quasi-variational inequality (QVI) (4.3) also admits solutions [chan_generalized_1982][Theorem 5.2]. However, as reported in [facchinei_generalized_2007, harker_generalized_1991, pang_quasi-variational_2005], solution methods for QVIs are scarce and typically require strong assumptions (for instance contractivity) on the problem structure. In both formulations, lower semicontinuity of the constraint maps remains a key requirement.
For each $x\in\mathcal{D}=\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{N}$, Assumption 3.1 ensures that ${\bf X}(x)$ is nonempty, convex, and compact. Combined with continuity and convexity of $\Psi(x,\cdot)$, this implies that $\Psi(x,\cdot)$ admits a minimizer on ${\bf X}(x)$. Hence, the merit function $\Phi:\mathcal{D}\to\mathbb{R}$,
$\Phi(x)=\min\limits_{y\in{\bf X}(x)}\Psi(x,y)$
is well defined. In some papers (see [hintermuller_generalized_2015, facchinei_generalized_2007]) the Nikaido–Isoda function and the merit function are defined as
$\tilde{\Psi}(x,y)=\Psi(x,x)-\Psi(x,y),\qquad\tilde{\Phi}(x)=\Psi(x,x)-\Phi(x),$
so that equilibria correspond to roots of $\tilde{\Phi}$. Here we keep Nikaido–Isoda’s original definition [nikaido_note_1955].
@@
When each ${\bf X}_{i}$ is given by (2.2) for some $\mathcal{C}\subset X$, the set of fixed points of $\bf X$} coincides with $\mathcal{C}$. This observation is important because every generalized Nash equilibrium must satisfy $\overline{x}\in{\bf X}(\overline{x})$, and therefore $\overline{x}\in\mathcal{C}$. In particular, in Lemma 4.1 and Lemma 4.2 the condition $\overline{x}\in{\bf X}(\overline{x})$ can be replaced by $\overline{x}\in\mathcal{C}$.
However, replacing all occurrences of ${\bf X}(\overline{x})$ by $\mathcal{C}$ would in general remove many equilibria, because ${\bf X}(\overline{x})$ depends on $\overline{x}$ while $\mathcal{C}$ does not. This effect is clearly visible in the two-company Cournot game of Example 2.3. For concreteness, let us fix $\eta=4$, $p=1$, $c_{1}=1$, $c_{2}=3/2$, and the shared capacity constraint
$x_{1}+x_{2}\leqslant C,$
so that
$\mathcal{C}=\{(x_{1},x_{2})\in\mathbb{R}_{+}^{2}:x_{1}+x_{2}\leqslant C\}.$
If $C\geqslant 11/6$, then the unique unconstrained Nash equilibrium lies in $\mathcal{C}$, so the generalized and the shared-constraint formulations coincide. If $C<11/6$, the intersection of the unconstrained best-response curves lies outside $\mathcal{C}$ and one obtains a whole segment of generalized Nash equilibria on the boundary $\{x_{1}+x_{2}=C\}$, as illustrated in Figure 2.
To see what happens when one replaces the pointwise feasible sets ${\bf X}_{i}(x_{-i})$ by the shared constraint $\mathcal{C}$ in the Nikaido–Isoda formulation, consider the case $C=1$. The associated Nikaido–Isoda function is
$\Psi(x,y)=J_{1}(y_{1},x_{2})+J_{2}(x_{1},y_{2})=-y_{1}\left(3-y_{1}-x_{2}\right)-y_{2}\left(\tfrac{5}{2}-x_{1}-y_{2}\right),$
where $x=(x_{1},x_{2})$ and $y=(y_{1},y_{2})$. If we now enforce the shared-constraint formulation by solving
$\min_{y\in\mathcal{C}}\Psi(\overline{x},y)\qquad\text{for $\overline{x}\in\mathcal{C}$,}$ (4.4)
then a direct Karush-Kuhn-Tucker (KKT) computation (or a variational-inequality argument) shows that
$\Psi(\overline{x},\overline{x})=\min\limits_{y\in\mathcal{C}}\Psi(\overline{x},y),\qquad\overline{x}\in\mathcal{C}$
holds if and only if $\overline{x}=(3/4,1/4)$. Thus, for $C=1$ the formulation in (4.4) yields one equilibrium, whereas with the original formulation (4.2), the game admits a continuum of equilibria along the boundary segment $x_{1}+x_{2}=1$.
Although, as this example shows, many equilibria are lost when ${\bf X}(\overline{x})$ is simply replaced by $\mathcal{C}$, the restricted notion is widely used because it simplifies analysis and computation and often provides a natural selection mechanism, which is crucial from a numerical viewpoint.
Equilibria obtained under this simplification are called variational equilibria.
@@
For simplicity of the statements we have implicitly assumed that $\mathcal{C}\subset X^{\rm{ad}}.$ If not, then we need to replace $\mathcal{C}$ with $X^{\rm{ad}}\cap\mathcal{C}$ in the theorem above.
@@
Obviously variational equilibria are Nash equilibria. The converse, as our example above shows, is not true and there are multiple other examples in the literature showing this fact, see for instance [facchinei_generalized_2007, harker_generalized_1991]. The most important games in which the constraint maps can be written as (2.2), are the jointly convex and potential games [hintermuller_generalized_2015, facchinei_generalized_2007, chan_generalized_1982].
Definition 4.6.
A jointly convex game is a game whose constraint maps are defined by (2.2) and $\mathcal{C}$ is, in addition to closed, convex.
Due to convexity, existence of variational equilibria follows, in our setting, if $\mathcal{C}$ is compact. For completeness we include the statement and sketch of the proof below.
Theorem 4.7.
Let $\mathcal{C}$ be a compact and convex subset of $X$ such that $X^{\rm{ad}}\cap\mathcal{C}\neq\emptyset$ and assume the structure (2.2) holds for the constraint maps. Then, every generalized game whose corresponding Nikaido–Isoda function $\Psi(\cdot,y)$ is upper semicontinuous for each $y\in\mathcal{C}$ has a variational equilibrium.
Proof.
It follows directly from the Ky-Fan theorem [iusem_new_2003, Theorem 1.1], which guarantees the existence of $\overline{x}\in\mathcal{C}$ such that
$\sup_{y\in\mathcal{C}}(\Psi(\overline{x},\overline{x})-\Psi(\overline{x},y))\leqslant 0.$ (4.7)
This completes the proof. ∎
In case $\mathcal{C}$ is not convex, existence can still be proved given more structure on the objective functions (or the preference maps). This leads to so-called potential games.
Definition 4.8 (Potential game).
Assume that the constraint maps are as in (2.2) and that $\mathcal{C}$ is closed. A generalized game is a potential game if there exists a continuous function $\mathcal{G}:X\to\mathbb{R}$, known as the potential function, such that for all $i$, all $x_{-i}$ and all $y_{i},z_{i}\in{\bf X}_{i}(x_{-i})$ the condition
$\mathcal{J}_{i}(y_{i},x_{-i})-\mathcal{J}_{i}(z_{i},x_{-i})>0$
implies
$\mathcal{G}(y_{i},x_{-i})-\mathcal{G}(z_{i},x_{-i})\geqslant g(\mathcal{J}_{i}(y_{i},x_{-i})-\mathcal{J}_{i}(z_{i},x_{-i})),$
where $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$, known as the forcing function, is such that $g(t_{k})\to 0$ implies $t_{k}\to 0.$
Theorem 4.9.
A potential game with $X^{\rm{ad}}$ compact and the forcing function $g$ continuous has an equilibrium.
Proof.
The proof is divided into two steps. First, consider the optimization problem
$\min_{x\in X^{\rm{ad}}\cap\mathcal{C}}\mathcal{G}(x).$ (4.8)
Since $X^{\rm{ad}}$ is compact and $\mathcal{C}$ is closed, it follows that $X^{\rm{ad}}\cap\mathcal{C}$ is compact. Since $\mathcal{G}$ is continuous, $\mathcal{G}(X^{\rm{ad}}\cap\mathcal{C})\subset\mathbb{R}$ is also compact. Hence there exists at least one minimizer $\hat{x}\in X^{\rm{ad}}\cap\mathcal{C}$ of $\mathcal{G}$. The second step is then to prove that if $\overline{x}$ is an optimal solution of (4.8), it is also a Nash equilibrium. Suppose not, then there exists at least one $i\in I$ and at least one $y_{i}\in{\bf X}_{i}(\overline{x}_{-i})$ such that $\mathcal{J}_{i}(\overline{x})>\mathcal{J}_{i}(y_{i},\overline{x}_{-i}).$ Hence, by definition of potential game we have
$\mathcal{G}(\overline{x})\geqslant\mathcal{G}(y_{i},\overline{x}_{-i})+g(\mathcal{J}_{i}(\overline{x})-\mathcal{J}_{i}(y_{i},\overline{x}_{-i})),$
and by properties of $g$ it follows that $\mathcal{G}(\overline{x})>\mathcal{G}(y_{i},\overline{x}_{-i}),$ a contradiction. ∎
Remark 4.10.
Observe that lower semicontinuity of $\mathcal{G}$ is enough for the theorem above to be true, although the proof would be slightly different.
Thus, finding a potential function $\mathcal{G}$ with a continuous forcing function $g$ is a sufficient condition to guarantee the existence of a Nash equilibrium in a potential game. As in the jointly convex case, the set of equilibria of a potential game is in general a proper subset of the full set of Nash equilibria. Every jointly convex game is, in particular, a potential game with $\mathcal{G}$ given by the Nikaido–Isoda function. The main existence results of this paper rely on the same philosophy: we replace strong regularity assumptions on the constraint maps by additional convexity or potential-type structure.
5 Potential lack of lower semicontinuity of constraint maps
In this section we show for convex games that lower semicontinuity of the constraint maps is not needed for equilibrium existence once additional geometric structure is available. This allows us to cover games with significantly more complex and heterogeneous constraints, such as generalized PDE-constrained games where each player’s state equation depends on forecasts of incomplete or partially available information, only.
Definition 5.1.
Our first theorem below shows that graph-convexity (in place of lower semicontinuity) is also sufficient for existence of equilibria.
Theorem 5.2.
${\rm Fix}({\bf X})\coloneqq\{x\in X^{\rm{ad}};x\in{\bf X}(x)\}$
is nonempty. Then the game $G=({\bf X},\mathcal{J})$ has an equilibrium.
It is important to emphasize here that graph-convexity does not replace lower semicontinuity as a weaker condition in general. In fact, it is well known that graph-convexity implies lower semicontinuity in the interior of the domain. A proof of this fact can be found in [rockafellar_variational_1998, Theorem 5.9(b)] in finite dimensions, and the same proof can be adapted to the Banach space case without much work. However, in infinite dimensions interiority is delicate, as we illustrate in the example below.
Example 5.3.
$C:=\left\{\,u\in L^{2}(0,1):\int_{0}^{1}u(x)\,dx=0\,\right\}.$
Since $C=\ker\ell$ for the continuous, surjective linear functional
$\ell(u):=\int_{0}^{1}u(x)\,dx,$
Define the closed convex set
$X_{i}^{\rm{ad}}=X_{i}=L^{2}(0,1),\qquad i=1,2.$
Finally, the player-specific constraint maps ${\bf X}_{i}:X_{-i}\to 2^{X_{i}}$ given by
${\bf X}_{i}(x_{-i}):=\begin{cases}K,&\displaystyle\int_{0}^{1}x_{-i}(x)\,dx=0,\\[3.0pt] \emptyset,&\displaystyle\int_{0}^{1}x_{-i}(x)\,dx\neq 0,\end{cases}$
${\bf X}(x_{1},x_{2})=\begin{cases}K_{1}\times K_{2},&x_{1}\in C,\ x_{2}\in C,\\[3.0pt] \emptyset,&\text{otherwise}.\end{cases}$
The graph of the joint constraint map is
${\bf Gr}({\bf X})=\left\{\,(x,y)\in X\times X:x_{1}\in C,\ x_{2}\in C,\ y_{1}\in K_{1},\ y_{2}\in K_{2}\,\right\}=(C\times C)\times(K_{1}\times K_{2}),$
${\rm dom}\ {\bf X}=\left\{\,x\in X:{\bf X}(x)\neq\emptyset\,\right\}=\left\{(x_{1},x_{2})\in X:x_{1}\in C,\ x_{2}\in C\right\}=C\times C,$
which has empty interior in $X$.
Lower semicontinuity of $\bf X$} fails at every $\overline{x}=(\overline{x}_{1},\overline{x}_{2})\in{\rm dom}{\bf X}$. Fix any $\overline{y}\in{\bf X}(\overline{x})=K_{1}\times K_{2}$. Since $C$ is a proper closed subspace of $L^{2}(0,1)$, we can choose sequences
This example is a genuine two-player generalized Nash game whose joint constraint map has a convex graph, but a domain with empty interior in the strong topology, and it illustrates how lower semicontinuity can fail in such “thin” situations. ∎
It follows from Lemma 4.1 that a bundle $\overline{x}$ is a GNE for a game if and only if
This characterization enables us to prove Theorem 5.2 in a rather elementary way; it extends the classical argument of Nikaido and Isoda [nikaido_note_1955] to the present setting.
Proof of Theorem 5.2.
Suppose that a GNE does not exist. By Lemma 4.1 and (5.1) this means, in particular, that for each $x\in{\rm Fix}({\bf X})$ there exists $y_{x}\in{\bf X}(x)\cap{\rm Fix}({\bf X})$ such that $\Psi(x,x)>\Psi(x,y_{x}).$ Fixing $y_{x}$, it follows from continuity of $\Psi$ that the set $\Psi_{x}\coloneqq\{z\in X:\,\Psi(z,z)>\Psi(z,y_{x})\}$ is nonempty (in fact it contains $x$) and open in $X$.
${\rm Fix}({\bf X})\subset\bigcup_{x\in{\rm Fix}({\bf X})}\Psi_{x}$
$g_{i}(x)=-\min\{\Psi(x,y_{x_{i}})-\Psi(x,x),0\}$
$G(x)=\dfrac{\sum_{i=1}^{r}g_{i}(x)y_{x_{i}}}{\sum_{i=1}^{r}g_{i}(x)}$
[…]
which is a contradiction. Hence a Nash equilibrium must exist. ∎
It is worth noting that in [hintermuller_generalized_2015, kanzow_multiplier-penalty_2019] the existence of variational Nash equilibria was established by using Kakutani’s fixed-point theorem and a K. Fan inequality, respectively. In this sense, the argument above provides an alternative route to existence of (variational) equilibria in the case of jointly convex games. From a structural viewpoint, the key difference is that here we work directly on the fixed-point set of the joint constraint map and exploit convexity of its graph, instead of building an auxiliary best-response map whose upper semicontinuity must be checked.
The example below, inspired by the viscosity-regularized spot market system of [hintermuller_generalized_2015, Section 5.2], illustrates how equilibrium existence based on graph-convexity applies even when the feasible sets are not jointly convex.
Example 5.4 (PDE-constrained spot market).
$y_{t}-\Delta y=\sum_{i=1}^{N}u_{i}\quad\text{in }Q,\qquad y=0\text{ on }\{0,1\}\times(0,T),\qquad y(\cdot,0)=0.$ (5.2)
The solution operator $S:X\to Y$ associated with (5.2) is linear and continuous. Note that above $L^{2}(\Omega)$, $H_{0}^{1}(\Omega)$ denote the usual Lebesgue and Sobolev spaces [MR2424078] and $L^{2}(0,T;L^{2}(\Omega))$ as well as $L^{2}(0,T;H_{0}^{1}(\Omega))$ are the associated Bochner spaces [MR895589].
Hence the product mapping ${\bf X}:={\bf X}_{1}\times\cdots\times{\bf X}_{N}$ is graph-convex on $X^{\rm{ad}}$. Hence, as long as ${\rm Fix}({\bf X})\neq\emptyset$ (which is a standard assumption to make) one can easily prove the existence of equilibrium as long as the objective functions satisfy Assumption 3.2. ∎
Definition 5.5.
Among the many properties of KKM maps (see for instance [beer_topologies_1993] and references therein) is the fact that $x\in F(x)$ for all $x\in X$, i.e., $F$ is nonempty-valued and all vectors in $X$ are fixed points.
Theorem 5.6.
Proof.
6 Multiplier bias and uniqueness of an equilibrium
The previous sections focused on existence. In many applications, however, the set of equilibria is large and may contain continua of points, as in Example 2.3. This section revisits Rosen’s diagonal strict convexity condition in our Banach space setting. The main message is twofold: (i) suitable structural conditions on a weighted pseudogradient single out a unique variational equilibrium, and (ii) changing the weights $r\in(\mathbb{R}_{+}^{\ast})^{N}$ can be interpreted as introducing a multiplier bias that selects one equilibrium among many. We keep the objective functionals fixed and bias only the way players react to marginal costs.
Let $G=({\bf X},\mathcal{J})$ be a jointly convex game in which the family of objective functionals satisfies Assumption 3.2. Consider the jointly convex game $G_{r}=({\bf X},r\cdot\mathcal{J})$ for some $r\in\mathbb{R}^{N}$. Here, $\cdot$’ is the componentwise multiplication of vectors. A well-known fact [rosen_existence_1965] is that if $r\in(\mathbb{R}_{+}^{*})^{N}$ (i.e., $r_{i}>0$ for all $i\in I$), then

### v2

@@ 3 Potential lack of lower semicontinuity of constraint maps
${\rm Fix}({\bf X})\coloneqq\{x\in X^{\rm{ad}};x\in{\bf X}(x)\}$
is nonempty. Then the game $G=({\bf X},\mathcal{J})$ has an equilibrium.
It is important to emphasize here that graph-convexity does not replace lower semicontinuity as a weaker condition in general. In fact, it is well known that graph-convexity implies lower semicontinuity in the interior of the domain. A proof of this fact can be found in [31, Theorem 5.9(b)] in finite dimensions, and the same proof can be adapted to the Banach space case without much work. However, in infinite dimensions interiority is delicate, as we illustrate in the example below.
Example 3.3.
$C:=\left\{\,u\in L^{2}(0,1):\int_{0}^{1}u(x)\,dx=0\,\right\}.$
Since $C=\ker\ell$ for the continuous, surjective linear functional
$\ell(u):=\int_{0}^{1}u(x)\,dx,$
Define the closed convex set
$X_{i}^{\rm{ad}}=X_{i}=L^{2}(0,1),\qquad i=1,2.$
Finally, the player-specific constraint maps ${\bf X}_{i}:X_{-i}\to 2^{X_{i}}$ are given by
${\bf X}_{i}(x_{-i}):=\begin{cases}K,&\text{if }\displaystyle\int_{0}^{1}x_{-i}(x)\,dx=0,\\[3.0pt] \emptyset,&\text{if }\displaystyle\int_{0}^{1}x_{-i}(x)\,dx\neq 0,\end{cases}$
${\bf X}(x_{1},x_{2})=\begin{cases}K_{1}\times K_{2},&\text{if }x_{1}\in C,\ x_{2}\in C,\\[3.0pt] \emptyset,&\text{otherwise}.\end{cases}$
The graph of the joint constraint map is
${\bf Gr}({\bf X})=(C\times C)\times(K_{1}\times K_{2}),$
${\rm dom}\ {\bf X}=\left\{\,x\in X:{\bf X}(x)\neq\emptyset\,\right\}=C\times C,$
which has empty interior in $X$.
Further, lower semicontinuity of $\bf X$} fails at every $\overline{x}\in{\rm dom}\ {\bf X}$. In fact, fix any $\overline{y}\in{\bf X}(\overline{x})=K_{1}\times K_{2}$. Since $C$ is a proper closed subspace of $L^{2}(0,1)$, we can choose sequences
The example above is a genuine two-player generalized Nash game whose joint constraint map has a convex graph, but a domain with empty interior in the strong topology, and it illustrates how lower semicontinuity can fail in such “thin” situations.
3.1 The Nikaido-Isoda Function
The Nikaido-Isoda function $\Psi:X\times X\to\mathbb{R}$ reads
@@ 3.1 The Nikaido-Isoda Function
Under Assumption 2.5, $\Psi$ is continuous and, for each $x\in X$, the map $\Psi(x,\cdot)$ is convex. A strategy bundle $\overline{x}\in X^{\rm{ad}}$ is a Nash equilibrium of the game $G=({\bf X},\mathcal{J})$ if and only if $\overline{x}\in{\bf X}(\overline{x})$ and $\Psi(\overline{x},\overline{x})\leqslant\Psi(\overline{x},y)$, for all $y\in{\bf X}(\overline{x})$. The forward implication of this equivalence is trivial. For the backwards one, assume that $\overline{x}$ is a fixed point of $\bf X$} and that $\Psi(\overline{x},\cdot)$ reaches its global minimum at $\overline{x}.$ If $\overline{x}$ is not a Nash equilibrium, then for some $i\in I$, there exists $y_{i}\in{\bf X}_{i}(\overline{x}_{-i})$ such that $\mathcal{J}_{i}(\overline{x})>\mathcal{J}_{i}(y_{i},\overline{x}_{-i})$. Noticing that $(y_{i},\overline{x}_{-i})\in{\bf X}(\overline{x})$, the result follows because
@@
In other words, $\overline{x}\in{\bf X}(\overline{x})$ is a GNE if and only if $\Psi(\overline{x},\overline{x})=\Phi(\overline{x})$ where $\Phi$ is defined as
$\Phi(x)=\min\left\{\Psi(x,y)\ \ {\rm s.t.}\ \ y\in{\bf X}(x)\right\}.$ (3.2)
@@
Among other consequences, this equivalence allows one to characterize Nash equilibria – under extra regularity of the objective functions – as the solution of a quasi-variational inequality (QVI). It also implies that a bundle $\overline{x}$ is a GNE for a game if and only if
We can then prove Theorem 3.2 in a rather elementary way; it extends the classical argument of Nikaido and Isoda [27] to the present setting.
3.2 Proof of Theorem 3.2
Suppose that a GNE does not exist. This means, in particular, that for each $x\in{\rm Fix}({\bf X})$ there exists $y_{x}\in{\bf X}(x)\cap{\rm Fix}({\bf X})$ such that $\Psi(x,x)>\Psi(x,y_{x}).$ Fixing $y_{x}$, it follows from continuity of $\Psi$ that the set $\Psi_{x}\coloneqq\{z\in X:\,\Psi(z,z)>\Psi(z,y_{x})\}$ is nonempty (in fact it contains $x$) and open in $X$.
${\rm Fix}({\bf X})\subset\bigcup_{x\in{\rm Fix}({\bf X})}\Psi_{x}$
$g_{i}(x)=-\min\{\Psi(x,y_{x_{i}})-\Psi(x,x),0\}$
$G(x)=\dfrac{\sum_{i=1}^{r}g_{i}(x)y_{x_{i}}}{\sum_{i=1}^{r}g_{i}(x)}$
[…]
which is a contradiction. Hence a Nash equilibrium must exist.
The example below, inspired by the viscosity-regularized spot market system of [19, Section 5.2], illustrates how equilibrium existence based on graph-convexity applies even when the feasible sets are not jointly convex.
Example 3.4 (PDE-constrained spot market).
$\begin{cases}y_{t}-\Delta y=\sum_{i=1}^{N}u_{i}\quad\text{in }Q,\\ y=0\text{ on }\{0,1\}\times(0,T),y(\cdot,0)=0.\end{cases}$ (3.4)
The solution operator $S:X\to Y$ associated with (3.4) is linear and continuous. Note that above $L^{2}(\Omega)$, $H_{0}^{1}(\Omega)$ denote the usual Lebesgue and Sobolev spaces [1] and $L^{2}(0,T;L^{2}(\Omega))$ as well as $L^{2}(0,T;H_{0}^{1}(\Omega))$ are the associated Bochner spaces [37].
Hence the product mapping ${\bf X}:={\bf X}_{1}\times\cdots\times{\bf X}_{N}$ is graph-convex on $X^{\rm{ad}}$. Hence, as long as ${\rm Fix}({\bf X})\neq\emptyset$ (which is a standard assumption to make) one can easily prove the existence of an equilibrium as long as the objective functions satisfy Assumption 2.5.
Definition 3.5.
Among the many properties of KKM maps (see for instance [4] and references therein) is the fact that $x\in F(x)$ for all $x\in X$, i.e., $F$ is nonempty-valued and all vectors in $X$ are fixed points.
Theorem 3.6.
Proof.
4 Multiplier bias and uniqueness of an equilibrium for jointly convex games
@@
When each ${\bf X}_{i}$ is given by (2.2) for some $\mathcal{C}\subset X$, the set of fixed points of $\bf X$} coincides with $\mathcal{C}$. This observation is important because every generalized Nash equilibrium must satisfy $\overline{x}\in{\bf X}(\overline{x})$, and therefore $\overline{x}\in\mathcal{C}$. In particular, if $\mathcal{C}$ is closed and convex, the corresponding game is called jointly convex. Jointly convex games allow for a notion of a subclass of Nash equilibria known as variational equilibria. This notion is known to be more restrictive - in the sense that many more Nash equilibria may exist - but more amenable to numerical computations. Hence, it is very important for applied game theory [11, 19, 10].
@@
For simplicity we have implicitly assumed that $\mathcal{C}\subset X^{\rm{ad}}$, otherwise we need to replace $\mathcal{C}$ with $X^{\rm{ad}}\cap\mathcal{C}$ in the theorem above.
@@
Obviously variational equilibria are Nash equilibria. The converse is not true and there are multiple examples in the literature showing this fact, see for instance [11, 16].
In many applications, the set of equilibria of a GNEP is large and may contain continua of points [11]. We want to end this section by extending Rosen’s diagonal strict convexity condition to our Banach space setting for jointly convex games. The main message is twofold: (i) suitable structural conditions on a weighted pseudogradient single out a unique variational equilibrium, and (ii) changing the weights $r\in(\mathbb{R}_{+}^{\ast})^{N}$ can be interpreted as introducing a multiplier bias that selects one equilibrium among many. We keep the objective functionals fixed and bias only the way players react to marginal costs.
Let $G=({\bf X},\mathcal{J})$ be a jointly convex game in which the family of objective functionals satisfies Assumption 2.5. Consider the jointly convex game $G_{r}=({\bf X},r\cdot\mathcal{J})$ for some $r\in\mathbb{R}^{N}$. Here, $\cdot$’ is the componentwise multiplication of vectors. A well-known fact [32] is that if $r\in(\mathbb{R}_{+}^{*})^{N}$ (i.e., $r_{i}>0$ for all $i\in I$), then